A circuit has a current of 1340 mA and a voltage of 11 V. What is the power of 

the circuit? What is the resistance of the circuit?


1
Expert's answer
2021-06-24T08:01:58-0400


Gives


Current (I)=1340mA

Voltage=11V

Power="IV"


"P=1340\times10^{-3}\times11=14.74watt"

Resistance


"R=\frac{P}{I^2}=\frac{14.74}{(1340\times10^{-3})^2}=8 .20\Omega"